OpenStudy (anonymous):

Find the values of a and b for the curve xy^2 + ay = bx such that at the point (1,-1), the curve has a slope of -1/2.

OpenStudy (anonymous):

You're familiar with derivatives, correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Well, we need to find the derivative of this function and then we need to go and plug in the in the value of x,y and set that= -1/2 a and b are considered to be constants in this case so can you differentiate this equation with respect to x and tell me what you get?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }= \frac{ b - y^2 }{ 2xy+a }\]

OpenStudy (anonymous):

OK, we have the derivative, correct. I'm trying to find a system of solutions to help find a and b,

OpenStudy (anonymous):

\[\frac{ b-y^2 }{ 2xy+a } = -\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

That's one equation. We need another one

OpenStudy (anonymous):

since y = -1 and x = 1. Shouldn't I switch the x's and y's to these numbers?

OpenStudy (mathstudent55):

Yes.

OpenStudy (anonymous):

You mean plug in 1 for x and -1 for y?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[xy^2 + ay = bx\]

OpenStudy (anonymous):

(1)(-1)^2 + a(-1) = b(1) 1 - a = b

OpenStudy (anonymous):

I'm not sure if that equation would work though. We can try

ganeshie8 (ganeshie8):

the given point must satisfy : f(x) = 0 f'(x) = -1/2

OpenStudy (mathstudent55):

Plug in (1, -1) both in the equation of the curve and in the equation of the derivative. That will give you two equations in a and b.

OpenStudy (anonymous):

so now I just solve the derivative of the function

OpenStudy (anonymous):

Yes what @mathstudent55 confidently corrected my thought lol

OpenStudy (anonymous):

\[\frac{ (1-a) - (-1)^2 }{ 2(1)(-1)+a } = -\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Ok and that simplifies?

OpenStudy (anonymous):

a = -2

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

oo I just plug in -2 to the 1 - a = b to find b

OpenStudy (anonymous):

1 - (-2) = b 1 + 2 = b b = 3 everyone left immediately =.= thanks though

OpenStudy (anonymous):

Correct!

OpenStudy (mathstudent55):

Excellent work, @Mateaus !

OpenStudy (anonymous):

Thanks everyone.

OpenStudy (mathstudent55):

You're welcome.